# theoTheorem

### v2

@@ III.3 Matrix element analysis
Further insight into the behavior of noise in the initial and final stages of the adiabatic state preparation is provided through the analysis of dominant matrix elements that facilitate phase noise excitation. Given that the Hamiltonian in Eq.(II) is defined in terms of the many-body number operator $\hat{n}=\sum_{k}^{N}\hat{n}_{k}$, it is more natural to work in the rotating frame, where phase noise acting on $\Omega$ manifests itself as a fluctuating detuning $\delta$ because it alters the laser frequency seen by the atom [40, 41]. This explains why laser phase noise fluctuations on Rabi frequency can be treated as a perturbation in detuning, with a new effective detuning term $\delta(t)\rightarrow\delta_{\mathrm{eff}}(t)=\delta(t)+\dot{\phi}(t)$ acting on the system through $\hat{n}$. Fig. 6 shows the available transitions that can be induced by phase noise, plotting all matrix elements of $\hat{n}$ for the four most prominent energy eigenstates in the initial and final Hamiltonian of the adiabatic state preparation. In the case of the initial Hamiltonian in Fig. 6(a), there is adifferential writesingle dominant transition from the ground state $|E_{0}\rangle$ to the ninth excited eigenstate $|E_{9}\rangle$ at an energy of $E_{9}=1.5V_{dd}$. As $|E_{9}\rangle$ becomes excited over time, its transition elements also become relevant, and we once again see the two dominant excitation channels are to excite to $|E_{64}\rangle$ with energy $3V_{dd}$, or de-excite back to $|E_{0}\rangle$. The pattern is repeated for higher eigenstates, such that time evolving a Hamiltonian with phase noise and a constant $\Omega/V_{dd}=1.0$ leads to the formation of an excitation ladder of equally spaced peaks (see Fig. 11(a) in Appendix B). However, the picture is very different in the final Hamiltonian of the adiabatic state preparation with $\Omega/V_{dd}=0.1$ in Fig. 6(b), where we observe excitation from the ground state as well as the higher energy states strongly suppressed. For $|E_{0}\rangle$, the most probable transition has a probability that is two orders of magnitude smaller than for the original Hamiltonian, suggesting that the ground state is much more robust to excitation in the final stages of the ramp. Furthermore, the energy eigenvectors that dominated the excitation dynamics in the initial Hamiltonian appear as higher excited states, while the equidistant transition peaks of the matrix elements of the initial Hamiltonian are replaced with transitions for much smaller probabilities and clustered near their respective eigenstate. This suggests that excitation due to laser phase noise is much more constrained in the final stages of the experiment, with smaller transition rates leading to less overall excitation.
[…]
@@ IV Thermalization
For systems that relax to a thermal equilibrium, the underlying mechanism is most robustly described by the eigenstate thermalization hypothesis (ETH) [29]. The ETH states that for nonintegrable systems (that do not exhibit many-body localization [53, 54]) the expectation values fluctuate negligibly around eigenstates $|\Psi_{\alpha}\rangle$ that are close to the energy of the system. Thus, each individual eigenstate can be thought of as an independent thermal state. In the thermodynamic limit, the resulting state should be constant and predictable by the total energy of the system. The hypothesis also presumes that the observable of interest is ‘well behaved’, in that $\langle\Psi_{\alpha}|\hat{A}|\Psi_{\alpha}\rangle$ behaves smoothly as a function of eigenstates of the Hamiltonian $|\Psi_{\alpha}\rangle$, with no sudden discontinuities in the vicinity of the expected energy that would create a sensitivity to slight changes in relaxation dynamics. Given these conditions the ETH predicts that the long time (LT) expectation value of a few body observable $\hat{A}$ with an eigenstate $|\Psi_{\alpha}\rangle$ of a many-body system Hamiltonian $\hat{H}$ with corresponding well defined total energy $E_{\alpha}$ is equal to the thermal ensemble $\langle\hat{A}\rangle_{\text{Therm}}(E_{\alpha})$ of $\hat{A}$ at a mean energy $E_{\alpha}$
$\langle\Psi_{\alpha}|\hat{A}|\Psi_{\alpha}\rangle_{LT}=\langle\hat{A}\rangle_{\text{Therm}}(E_{\alpha}).$ (5)
Hence, the ETH provides a framework for comparing expectation values which are directly dependent on initial conditions to thermal ensembles which instead depend only on the total energy of the system. This implies a universal equilibrium for a thermalizing system across many trajectories of equivalent energy.
It will be necessary to define what is meant by a long time average $\langle\Psi_{\alpha}|\hat{A}|\Psi_{\alpha}\rangle_{LT}$ in Eq. (5). Any dynamics in a quantum system leads to correlations between excited energy eigenstates that cause fluctuations in measured expectation values as a function of time and energy difference. The time it takes for these fluctuations to die down drops with increasing system size, but for small and intermediate systems with large energy gaps this time can be too long to measure or even simulate. However, numerically a long time average can be accessed directly by considering the diagonal ensemble
$\langle\hat{A}\rangle_{LT}=\sum_{i}|c_{ii}|^{2}\langle E_{i}|\hat{A}|E_{i}\rangle=\sum_{i}|c_{ii}|^{2}\hat{A}_{ii},$ (6)
@@
IV.1 Evaluating thermalization
@@
where $|\psi\rangle$ represents the state of the system at the end of the adiabatic state preparation such that the left hand side of the equation gives the final energy at the end of the procedure. Once a value for $\beta$ is determined it can be used in Eq. (7) to evaluate arbitrary thermal expectation values at the appropriate energy. Given that phase noise has the potential for exciting across the full energy spectrum it is important to access the full Hilbert space, meaning that we limit ourselves to a system size of 11 sites which can be completely diagonalized.
@@
Fig. 7 provides all the distributions required to generate long time and thermal expectation values as a function of energy relative to the ground state. Figures 7(a) and 7(b) show the two observables chosen for the comparison to be the interaction energy $\hat{H}_{int}$ as defined in Eq. (2), as well as a $\hat{O}_{Z_{2}}$ order parameter defined in Eq. (3) that measures long range $Z_{2}$ ordering that is maximally represented in the ground state. Given the relatively small Rabi frequency $\Omega=0.1V_{dd}$ at the end of the state preparation, we are investigating a Hamiltonian that is highly integrable, with structure in energy levels leading to regions of clustered eigenvalues (see Fig. 10 in Appendix B). Although neither of these observables vary smoothly as a function of energy as prescribed by the ETH, there is considerably less variance in $H_{int}$ compared to the $\hat{O}_{Z_{2}}$ order parameter meaning that small changes in excitation energy should affect expectation values of the latter more. The distributions as a function of eigenstate energy for $H_{int}$ and the $\hat{O}_{Z_{2}}$ order parameter are then weighted by the diagonal and canonical ensembles. Fig. 7(c) provides example distributions for a short preparation of $T_{3}V_{dd}=45$ and a much longer preparation at $T_{3}V_{dd}=400$ in which the Boltzmann distribution skews visibly towards higher excited states. The diagonal ensemble for the shorter $T_{3}V_{dd}=45$ shows the first excited state having the highest representation, with a majority of higher energy states left largely unexcited and thus with an lower occupation than is predicted by the Boltzmann distribution. For a much longer preparation of $T_{3}V_{dd}=400$ the first excited state drops far below the occupation of the Boltzmann distribution, as diabatic excitation is subdued. Instead, we see a general rise in occupancy across the entire energy spectrum, with a cohort of low energy eigenstates clustered below $E=1V_{dd}$ from the ground state being most prominent.
@@
Figure 7: (a) Interaction energy $H_{int}$ defined in Eq. (2) and (b) $\hat{O}_{Z_{2}}$ ordering parameter as a function of low energy eigenstates of the target Hamiltonian with $\Omega=0.1V_{dd}$ and $\delta=1.1V_{dd}$ in a one dimensional Rydberg spin chain with 11 sites. (c) Long time ensemble (LT) $\langle\Psi|E_{i}\rangle$, and thermal ensemble (Th) as predicted by the canonical ensemble from Eq. (7), for the final step of the adiabatic state preparation with durations $\text{T}_{3}=45,400tV_{dd}$. Energy defined with respect to ground state energy.
@@
Averaging over the operator distributions of Figures 7(a) and 7(b) weighted by the long time and thermal ensembles of a particular $T_{3}$ result in a single expectation value. Fig. 8 provides a comparison between thermal and long time expectation values of $\hat{H}_{int}$ and the $\hat{O}_{Z_{2}}$ order parameter, across a chain of 11 sites, and for a variety of $T_{3}$. The general relationship between both types of expectation value and time $T_{3}$ is roughly consistent with the results in Fig. 4, as the optimal ramp time coincides with the lowest interaction energy in the system and therefore the highest fidelity with the $Z_{2}$ ordered ground state. For durations shorter than the optimal fidelity, where diabatic excitation dominates, we see a clear divergence between the long time and thermal expectation values. The cause of this divergence can be seen in Fig. 7(c), which shows that for the shorter preparation time of $T_{3}=45tV_{dd}$ a majority of energy eigenstates have occupation probabilities below the Boltzmann distribution. However, as the duration of $T_{3}$ passes its optimum, and laser phase noise begins to dominate excitation, we see a clear convergence of the long time and thermal expectation values. In the case of $H_{int}$ the convergence is strong as soon as $T_{3}V_{dd}=60$, coinciding with a plateau of minimum energy seen for the same system size of 11 sites in Fig. 4(b). At the optimal $T_{3}$ there is not enough energy added to the system to differentiate the two expectation values. However, this strong convergence persists as phase noise excitations heat the system. In the case of the $\hat{O}_{Z_{2}}$ order parameter, a similar divergence of the two expectation values occurs at short $T_{3}$, along with convergence around the optimal time. For longer $T_{3}$, the expectation values for this parameter behave more erratically as can be expected given the large variance of the operator as a function of energy. Despite this, there is still good convergence between the thermal and long time expectation values even for the $\hat{O}_{Z_{2}}$ operator.
@@
Analysis of the diagonal ensemble for a ground state evolution of constant Hamiltonians at different $\Omega$ along the simulated adiabatic preparation reveals that phase noise excitation is considerably more prominent in the initial stages of the ramp (see Fig. 11 in Appendix B). When $\Omega$ is large enough to break integrability, and the phase noise is given enough time, the highly delocalized transitions that dominate excitation form energy peaks across the entire spectrum that remain visible even at the end of the simulation in Fig. 7(c). However, by the end of the state preparation when $\Omega$ decreases and the Hamiltonian becomes integrable, these excitation peaks become distorted as energy eigenstates shuffle into tightly packed clusters separated by large energy gaps. There is an overall drop in transition probability as highly delocalized excitations disappear, limiting the energy mobility around the system to the local heating of neighboring eigenstates predicted by Fig. 6(b) and shown in Fig. 11(c). The combination of aggressive delocalized excitation that facilitates rapid energy transfer in the first half of the ramp, and the slow localized heating of this transferred energy in the latter half seems to excite the system in a way such that averaged behavior is lockstep with the uniform predictions of the Boltzmann distribution, albeit the distributions differ considerably. In this way, we can say the adiabatic state preparation facilitates the convergence of the long-time and thermal ensemble. However, when phase noise is left uncontrolled in a non-integrable regime it leads to a diagonal ensemble that rises above the canonical ensemble and does not resemble thermal equilibrium. Similarly, evolving in a Hamiltonian that is fully or close to integrable leads to much less total energy transfer, with strictly localized excitation inhibiting the uniform excitation of the full energy spectrum.

### v3

@@ III.3 Matrix element analysis
Figure 6: Matrix elements of the many-body operator $\hat{\sigma}_{z}=\sum_{k}^{N}\hat{\sigma}^{k}_{z}$ for Hamiltonians occurring at different stages of the adiabatic state preparation from Figure 1 for an 11 site spin chain. Hamiltonian (a), with laser parameters $\Omega/\text{V}_{dd}=0.6$ and $\delta/\text{V}_{dd}=-1.0$ occurs in the middle of stage 2 of the protocol. Hamiltonian (b), with parameters $\Omega/\text{V}_{dd}=0.6$ and $\delta/\text{V}_{dd}=1.1$, occuring at the transition point between stages 2 and 3. Hamiltonian (c), with parameters $\Omega/\text{V}_{dd}=0.1$ and $\delta/\text{V}_{dd}=1.1$, occurring at the end of stage 3. Dashed lines mark energies of the ground state (blue) along with the energy of the state with the highest ground state transition rate (yellow), and subsequent states corresponding to the highest transition rates from the preceding state (green,red), while dots of matching color represent off-diagonal matrix elements.
Further insight into the behavior of noise in the initial and final stages of the adiabatic state preparation is provided through the analysis of dominant matrix elements that facilitate phase noise excitation. Since the Hamiltonian in Eq.(II) is defined in terms of the many-body number operator $\hat{n}=\sum_{k}^{N}\hat{n}_{k}$, it is useful to work in the rotating frame of the atoms themselves, which see the phase factor in the transverse drive as a stochastic fluctuation in detuning $\delta(t)\rightarrow\delta(t)-\dot{\phi}(t)$, where $\dot{\phi}(t)$ is the instantaneous change in frequency due to noise [26, 19]. We achieve this change of frames by applying the unitary transformation $\hat{U}_{t}=\mathrm{exp}({-i\phi(t)}\sum\hat{n}_{k})$, as shown in Appendix B. Figure 6 shows the distribution of transition rates between different many-body energy eigenstates for the Hamiltonian defined in Eq. (II) with parameters that occur at three different points throughout stages 2 and 3 of the adiabatic state preparation studied. Each plot shows transitions for the ground state $|E_{0}\rangle$, along with transitions for the state with the highest transition rate with $|E_{0}\rangle$ and again with states that have the highest transition rates with the preceding analyzed state. In this way, we trace the most probable excitation path for a population starting in the ground state, getting insight on how excitation dynamics changes throughout the protocol. We begin in Figure 6(a) with the Hamiltonian occurring in the middle of stage 2, where the dominant transition is that from the ground state $|E_{0}\rangle$ to the sixth excited eigenstate $|E_{6}\rangle$ at an energy of $E_{6}/V_{dd}~\approx 1.3$. The remaining cohort of non-negligible transition probabilities is at least an order of magnitude smaller, found at energies similar to $E_{6}$ and to a lesser degree around $E/V_{dd}~\approx 3.3$. As $|E_{6}\rangle$ becomes excited over time, its transition elements also become relevant, with two dominant excitation channels that excite upward to $|E_{31}\rangle$ with energy $E_{31}/V_{dd}\approx 2.7$ and other nearby states, or de-excite back to $|E_{0}\rangle$ leading to an oscillatory behavior between the two states. This behavior is repeated for $|E_{31}\rangle$. The increase in detuning throughout stage 2 not only decreases the energy gaps between states, making diabatic excitation more likely, but also increases the statistical weight across all off-diagonal elements. Moreover, a high transverse field of $\Omega/V_{dd}=0.6$ induces state mixing, which promotes higher energy transitions as the system in a more non-integrable regime (a discussion on integrability in this system is provided in Appendix C.1). Consequently, Stage 2 reconfigures the energetic structure of the many body system to facilitate not only initial phase noise excitation from the ground state, but also subsequent excitations that disperse energy across the entire spectrum. The culmination of these effects can be seen at the transition point between stages 2 and 3 in Figure 6(b) with laser parameters $\Omega/V_{dd}=0.6$ and $\delta/V_{dd}=1.1$, where we see a dramatic increase in both the density of available states and the magnitude of their respective transition rates. The highlighted excitation path of $|E_{0}\rangle\rightarrow|E_{13}\rangle\rightarrow|E_{88}\rangle\rightarrow|E_{288}\rangle$ is the most dominant, forming an array of equidistant peaks that can allow fast transfer of energy from the ground state to higher energy sectors. Furthermore, the high number of eigenstates with an elevated transition rate give phase noise further outlet to excite the system at a wider range of frequencies than at any other point throughout the adiabatic process. Transitions for the final Hamiltonian of the protocol are shown in Figure 6(c), with $\Omega/V_{dd}~=0.1$. Here we see excitation probability from the ground state as well as the higher energy states become strongly suppressed. For $|E_{0}\rangle$, the most dominant transition has a probability that is two orders of magnitude smaller than the one in Figure 6(b), suggesting that the ground state is much more robust against noise excitation in the final stages of the ramp. Moreover, the energy eigenvectors that dominated the excitation dynamics in Figure 6(b) appear as higher excited states, and the equidistant transition peaks of the matrix elements are replaced with transitions for much smaller probabilities that cluster near their respective eigenstate. This suggests that excitation due to laser phase noise is much more constrained in the final stages of the experiment, with available excitation states having smaller transition rates that are all near in energy.
The differences between the last two sets of matrix elements can be explained by the competition between the two laser parameters in the Hamiltonian in Eq. (II). As $\Omega/V_{dd}$ decreases, the integrability of the Hamiltonian changes, which has strong implications on the transition element spectrum. This change transitions the system from a non-integrable state with large state mixing to an integrable state that approaches the Ising model [45, 15]. Higher integrability leads to an underlying structure in the energy levels, which form clusters separated by large energy gaps (see Figure 11 in Appendix C.2). Such energy gaps, coupled with the absence of matrix elements with distant spectral indices that could facilitate excitation to different energy regions, inhibit the spread of excitation energy throughout the Hilbert space. Of course, the picture is much more complicated in the dynamical setting of the adiabatic state preparation where a time dependent $\Omega(t)$ constantly changes excitation patterns, but Figure 6 along with phase noise evolutions of time-independent Hamiltonians at various stages of the preparation (see Figure 12 in Appendix C.3) provide evidence that phase noise excitation is highly interlinked with the integrability of the system. As the transverse field that breaks integrability decreases, so does the ability of the phase noise to excite the system.
@@ IV Thermalization
For systems that relax to a thermal equilibrium, the underlying mechanism is most robustly described by the eigenstate thermalization hypothesis (ETH) [15]. The ETH states that for nonintegrable systems (that do not exhibit many-body localization [1, 20]) the expectation values fluctuate negligibly around eigenstates $|E_{i}\rangle$ that are close to the energy of the system. Thus, each individual eigenstate can be thought of as an independent thermal state. In the thermodynamic limit, the resulting state should be constant and predictable by the total energy of the system. The hypothesis also presumes that the observable of interest is ‘well behaved’, in that $\langle E_{i}|\hat{A}|E_{i}\rangle$ behaves smoothly as a function of eigenstates of the Hamiltonian $|E_{i}\rangle$, with no sudden discontinuities in the vicinity of the expected energy that would create a sensitivity to slight changes in relaxation dynamics. Given these conditions the ETH predicts that the long time (LT) expectation value of a few body observable $\hat{A}$ with an eigenstate $|E_{i}\rangle$ of a many-body system Hamiltonian $\hat{H}$ with a corresponding well defined energy $E_{i}$ is equal to the thermal ensemble $\langle\hat{A}\rangle_{\text{Therm}}(E_{i})$ of $\hat{A}$ at a mean energy $E_{i}$
$\langle E_{i}|\hat{A}|E_{i}\rangle_{LT}=\langle\hat{A}\rangle_{\text{Therm}}(E_{i}).$ (5)
Hence, the ETH provides a framework for comparing expectation values which are directly dependent on initial conditions to thermal ensembles which instead depend only on the total energy of the system. This implies a universal equilibrium for a thermalizing system across many trajectories of equivalent energy. It is then necessary to define exactly what is meant by a long time average $\langle E_{i}|\hat{A}|E_{i}\rangle_{LT}$ in Eq. (5). Any dynamics in a quantum system lead to a quantum state $|\psi\rangle=\sum_{i}c_{i}|E_{i}\rangle$ with correlations $c^{*}_{i}c_{j}$ between different excited states $|E_{i}\rangle$ and $|E_{j}\rangle$ that cause fluctuations in measured expectation values as a function of time and energy difference
$\langle\psi(t)|\hat{A}|\psi(t)\rangle=\sum_{i,j}e^{-\frac{i}{\hbar}(E_{i}-E_{j})t}c^{*}_{i}c_{j}\langle E_{i}|\hat{A}|E_{j}\rangle.$ (6)
The time it takes for these fluctuations to die down drops with increasing system size, but for small and intermediate systems with large energy gaps this time can be too long to measure or even simulate. However, numerically a long time average can be accessed directly by considering the diagonal ensemble
$\langle\psi|\hat{A}|\psi\rangle_{LT}=\sum_{i}|c_{ii}|^{2}\langle E_{i}|\hat{A}|E_{i}\rangle=\sum_{i}|c_{ii}|^{2}\hat{A}_{ii},$ (7)
@@
IV.1 Changing parameter regimes
Figure 7: (a) Plot of fidelity against decreasing values of $\Omega/V_{dd}$ for a modified version of the stage 3 adiabatic ramp discussed in Figure 1, now starting from $\Omega/V_{dd}=1.0$. Simulations are performed in an 11 site system with no phase noise, as colored lines show the fidelity for different evolution times with faster ramps incurring more diabatic excitation. (b,c) Overlap of final state and energy eigenstates after a ground state evolution under a time independent Hamiltonian from Eq. (II) for a total time of $TV_{dd}=400$ with added phase noise. Plots show simulation for different laser parameters occurring in the modified stage 3 ramp with, detuning set to $\delta/V_{dd}=1.1$ and Rabi frequency set to (b) $\Omega/V_{dd}=1.0$ and (c) $\Omega/V_{dd}=0.1$. Red triangles mark the ground state fidelity, $\mathrm{F}_{E_{0}}$, which is also provided in the legend. Error bars plot standard error across 100 time evolutions with independent laser phase noise signals using the same noise profile from Figure 3, but are too small to be seen on the log scale suggesting consistent excitation.
To better judge the thermalizing effect of phase noise, it is useful to shift our simulations to a regime in which noise excitation is stronger. In the previous sections, the focus was placed on studying the impact of laser phase noise on high-fidelity adiabatic state preparation. Operating in the $V_{dd}/2\pi=10~\mathrm{MHz}$ regime kept the characteristic system frequencies well above the dominant noise peak near $0.48~\mathrm{MHz}$ (see Figure 3), which suppressed excitation and maintained the system close to its instantaneous ground state. In contrast, this section examines the dynamics from the opposite perspective; rather than mitigating phase noise, we exploit its coupling to the system to study how noise can drive energy redistribution and apparent thermalization in an interacting Rydberg chain. For this purpose, we shift to a lower frequency interaction regime of $V_{dd}/2\pi=0.16~\mathrm{MHz}$ that is also studied in Figure 12 in Appendix C.3, where the characteristic transition energies of the Hamiltonian become comparable to the spectral components of the laser noise. In this new regime, preparation of high-fidelity adiabatic states is no longer feasible, as shown in Figure 5(b), in which the fidelity of the prepared state decreases to below 20% even for the smallest system size of 11 sites. Instead, our analysis of thermalization is isolated to stage 3 of the adiabatic ramp, in which phase noise dynamics undergo a dramatic change as the system moves from high to low integrability when the transverse field is reduced. To provide more time for the noise, we prolonge this transition by starting the adiabatic ramp from $\Omega/V_{dd}=1.0$ as opposed to $\Omega/V_{dd}=0.6$ as was the case in Section III.
Figure 7 demonstrates the key changes in the dynamics of the modified adiabatic ramp for stage 3. Firstly, Figure 7(a) plots fidelity with instantaneous ground state against the full range of $\Omega/V_{dd}$ for simulations of an 11 site system with no phase noise and for a variety of evolution times. We begin by initializing in the ground state of the initial Hamiltonian of our modified ramp with laser parameters $\Omega/V_{dd}=1.0$ and $\delta/V_{dd}=1.1$. Simulations without laser phase noise reveal that the main loss of fidelity due to diabatic excitation is now shifted to the middle of stage 3, coinciding with the critical region in the ground state energy gap that occurs at $\Omega/V_{dd}\approx 0.55$ (see Figure 1(c). Following this, Figures 7(b,c) provide the energy eigenstate overlap of ground state evolved with phase noise under two time independent Hamiltonians that occur at the beginning and end of the adiabatic protocol. We evolve an 11 site system for a long evolution time of $TV_{dd}=400$ to allow enough time for excitation, and observe a stark difference in the amount and nature of phase noise excitation in the two Hamiltonians. Simulations in Figure 7(b) at $\Omega/V_{dd}=1.0$ result in considerable excitation and a final ground state fidelity of 55.055.0%. Moreover, we observe a distinct excitation profile of equidistant peaks corresponding to states with dominant transition rates and high separation in the energy spectrum (see Figure 6(b) for the same mechanism in $\Omega/V_{dd}=0.6$), which allow a rapid transfer of energy from the ground state to higher energy sectors. However, below $\Omega/V_{dd}=0.6$ the energy landscape begins to undergo dramatic changes as energy states begin to cluster together, and by the time we reach the final Hamiltonian in Figure 7(c) the transition rates are suppressed, especially for eigenstates with large energy gaps (see Figure 11 in Appendix C.2). As a consequence, the final ground state fidelity of time independent Hamiltonian with $\Omega/V_{dd}=0.1$ remains at $90.3\%$, as phase noise finds it much harder to excite the system.
IV.2 Evaluating thermalization
@@
where $|\psi\rangle$ represents the state of the system at the end of the adiabatic state preparation such that the left-hand side of the equation gives the final energy at the end of the procedure averaged over 100 unique realizations. Once a value for $\beta$ is determined it can be used in Eq. (8) to evaluate arbitrary thermal expectation values at the appropriate energy. Given that phase noise has the potential for exciting across the full energy spectrum it is important to access the full Hilbert space, meaning that we limit ourselves to a system size of 11 sites which can be completely diagonalized.
@@
Figure 8 provides all the distributions required to generate long time and thermal expectation values as a function of energy relative to the ground state. Figures 8(a) and 8(b) show the two observables chosen for the comparison to be the interaction energy $\hat{H}_{int}$ as defined in Eq. (2), as well as a $\hat{O}_{Z_{2}}$ order parameter defined in Eq. (3) that measures long range $Z_{2}$ ordering that is maximally represented in the ground state. Given the relatively small Rabi frequency $\Omega=0.1V_{dd}$ at the end of the state preparation, we are investigating a Hamiltonian that is highly integrable, with structure in energy levels leading to regions of clustered eigenvalues (see Figure 11 in Appendix C). Although neither of these observables vary smoothly as a function of energy as prescribed by the ETH, there is considerably less variance in $H_{int}$ compared to the $\hat{O}_{Z_{2}}$ order parameter meaning that small changes in excitation energy should affect expectation values of the latter more. The distributions as a function of eigenstate energy for $H_{int}$ and the $\hat{O}_{Z_{2}}$ order parameter are then weighted by the diagonal and canonical ensembles. Figure8(c) provides example distributions for a short preparation of $T_{3}V_{dd}=45$ and a much longer preparation at $T_{3}V_{dd}=400$. Practically speaking, the determining factors of thermalization will be: how well the uniform Boltzmann distribution of the canonical ensemble (with an equivalent many-body energy) acts as a ‘line of best fit’ for the long-time distribution, and, at energies where these distributions disagree, how does the local behavior of the tested observable exacerbate or lessen the impact on the averaged expectation value. Moreover, since we are dealing with ground sate preparation, most of the energy will be held in the low-energy spectrum even after longer evolution times, meaning that the low-energy eigenvalues will have an out-sized impact on the final expectation values. The diagonal ensemble for the shorter $T_{3}V_{dd}=45$ shows the first excited state having the highest representation, with a majority of higher energy states left largely unexcited and thus with an lower occupation than is predicted by the Boltzmann distribution. For a much longer preparation of $T_{3}V_{dd}=400$ the first excited state drops far below the occupation of the Boltzmann distribution, as diabatic excitation is subdued. Instead, we see a general rise in occupancy across the entire energy spectrum, with a cohort of low energy eigenstates clustered below $E=1V_{dd}$ from the ground state being most prominent.
@@
Figure 8: (a) Interaction energy $H_{int}$ defined in Eq. (2) and (b) $\hat{O}_{Z_{2}}$ ordering parameter as a function of low energy eigenstates of the target Hamiltonian with $\Omega=0.1V_{dd}$ and $\delta=1.1V_{dd}$ in a one dimensional Rydberg spin chain with 11 sites. (c) Long time ensemble (LT) $\langle\Psi|E_{i}\rangle$, and thermal ensemble (Th) as predicted by the canonical ensemble from Eq. (8), for the final step of the adiabatic state preparation with durations $\text{T}_{3}V_{dd}=45,400$. Energy is defined with respect to $E_{gr}$, and error bars show standard error across 100 noise realizations.
@@
Averaging over the operator distributions of Figures 8(a) and 8(b) weighted by the long time and thermal ensembles of a particular $T_{3}$ result in two separate expectation values. Figure 9 provides a comparison between thermal and long time expectation values of $\hat{H}_{int}$ and the order parameter $\hat{O}_{Z_{2}}$, across a chain of 11 sites, and for a variety of $T_{3}$. Despite the regime change, the general relationship between both types of expectation value and time $T_{3}$ is roughly consistent with the results in Figure 4, as the optimal ramp time when minimal energy is added to the system coincides with the lowest interaction energy in the system and therefore the highest fidelity with the $Z_{2}$ ordered ground state. For durations shorter than the optimal fidelity, where diabatic excitation dominates, we see a clear divergence between the long time and thermal expectation values. The cause of this divergence can be seen in Figure 8(c), which shows that for the shorter preparation time of $T_{3}V_{dd}=45$ a majority of energy eigenstates have occupation probabilities below the Boltzmann distribution. However, as the duration of $T_{3}$ passes its optimum, and laser phase noise begins to dominate excitation, we see a clear convergence of the long time and thermal expectation values. In the case of $H_{int}$ the convergence is strong as soon as $T_{3}V_{dd}=60$, coinciding with a plateau of minimum energy seen for the same system size of 11 sites in Figure 4(b). At the optimal $T_{3}$ there is not enough energy added to the system to differentiate the two expectation values. However, this strong convergence persists as phase noise excitations heat the system. In the case of the $\hat{O}_{Z_{2}}$ order parameter, a similar divergence of the two expectation values occurs at short $T_{3}$, along with convergence around the optimal time. For longer $T_{3}$, the expectation values for this parameter behave more erratically as can be expected given the large variance of the operator as a function of energy. Moreover, the long time expectation value is consistently above the thermal prediction as energy in the long time ensemble is spread over only the ground state and a few energy sectors with high transition rates, meaning that there is always some more occupancy in the perfectly $Z_{2}$ ordered ground state than would be predicted by the Boltzmann distribution that uniform exponential decay across all eigenstates. Despite this, there is still good convergence between the thermal and long time expectation values even for the $\hat{O}_{Z_{2}}$ operator.
@@
Analysis of the diagonal ensemble for a ground state evolution of constant Hamiltonians at different $\Omega$ along the simulated adiabatic preparation reveals that phase noise excitation is considerably more prominent in the initial stages of the ramp (see Figure 12 in Appendix C.3). When $\Omega$ is large enough to break integrability, and the phase noise is given enough time, the highly delocalized transitions that dominate excitation form energy peaks across the entire spectrum that remain visible even at the end of the simulation in Figure 8(c). However, by the end of the state preparation when $\Omega$ decreases and the Hamiltonian becomes integrable, these excitation peaks become distorted as energy eigenstates shuffle into tightly packed clusters separated by large energy gaps. There is an overall drop in transition probability as highly delocalized excitations disappear, limiting the energy mobility around the system to the local heating of neighboring eigenstates predicted by Figure 6(b) and shown in Figure 12(d). The combination of aggressive delocalized excitation that facilitates rapid energy transfer in the first half of the ramp, and the slow localized heating of this transferred energy in the latter half seems to excite the system in a way such that averaged behavior is lockstep with the uniform predictions of the Boltzmann distribution, albeit the distributions differ considerably. In this way, we can say the adiabatic state preparation facilitates the convergence of the long-time and thermal ensemble. However, when phase noise is left uncontrolled in a non-integrable regime it leads to a diagonal ensemble that rises above the canonical ensemble and does not resemble thermal equilibrium. Similarly, evolving in a Hamiltonian that is fully or close to integrable leads to much less total energy transfer, with strictly localized excitation inhibiting the uniform excitation of the full energy spectrum.